OpenStudy (anonymous):

determine if the integral is convergent or divergent \[\int\limits_{0}^{\infty} 1/\sqrt{x+1}\]

OpenStudy (experimentx):

looks like diverges.

OpenStudy (turingtest):

probably, yeah but how to show it...

OpenStudy (anonymous):

i don't know hot to show it.

OpenStudy (turingtest):

oh I know, comparison

OpenStudy (anonymous):

i need to learn these for my test so i can't depend on wolfram

OpenStudy (anonymous):

ways to show it :)

OpenStudy (experimentx):

integrate it --> that will give infinity.

OpenStudy (turingtest):

oh yeah, duh... :P

OpenStudy (experimentx):

also summation of 1/sqrt(x+1) <--- itself seems to diverge <-- by ratio test.

OpenStudy (experimentx):

@TuringTest done something like this before?? http://openstudy.com/study#/updates/4f883247e4b0505bf08753ec

OpenStudy (anonymous):

i'm so confused.

OpenStudy (anonymous):

i'll try to integrate it now.

OpenStudy (anonymous):

2\[2\sqrt{\infty+1}-2\sqrt{0+1}\]

OpenStudy (turingtest):

in general for indefinite integrals we want to write\[\int_{0}^{n}\frac1{\sqrt{1+x}}dx\]then take the limit as \(n\to\infty\)

OpenStudy (turingtest):

@azalea that's a very unconventional way to write it. Your teacher may not like that.

OpenStudy (turingtest):

also it's not quite right...

OpenStudy (anonymous):

LOL. oh ok

OpenStudy (experimentx):

I thought you were doing integral test for convergence of 1/sqrt(1+x)

OpenStudy (turingtest):

try what I said\[\int_{0}^{n}\frac1{\sqrt{1+x}}dx\]then take the limit as n goes to infinity

OpenStudy (experimentx):

I don't like it too.

OpenStudy (anonymous):

yeah i did it wrong

OpenStudy (anonymous):

\[\lim_{0 \rightarrow t}\int\limits_{0}^{t}1/\sqrt{(x+1)}\]

OpenStudy (anonymous):

where t=infinity

OpenStudy (experimentx):

dx missing.

OpenStudy (turingtest):

not quite... limit is wrong too

OpenStudy (anonymous):

oh yeah. i forgot the dx.

OpenStudy (turingtest):

\[\lim_{n\to\infty}\int_{0}^{n}\frac1{\sqrt{1+x}}dx\]

OpenStudy (anonymous):

i meant t--> infinity

OpenStudy (anonymous):

\[\lim_{t \rightarrow \infty}2\sqrt{(x+1)} \]

OpenStudy (anonymous):

then as the limit approaches infinity it's infinity

OpenStudy (turingtest):

you should have already plugged in the t, and you forgot the evaluation at 0, but yeah

OpenStudy (anonymous):

oh yeah i didn' finish

OpenStudy (anonymous):

\[2\sqrt{(t+1)}-2\sqrt{(0+1)}\]

OpenStudy (turingtest):

yup, you got it :)

OpenStudy (anonymous):

\[2\sqrt{(t+1)}-2\]

OpenStudy (experimentx):

put the limit there.